# Theoretical Insights into Overparameterized Models in Multi-Task and Replay-Based Continual Learning

### v1

@@ Abstract
Multi-task learning (MTL) is a machine learning paradigm that aims to improve the generalization performance of a model on multiple related tasks by training it simultaneously on those tasks. Unlike MTL, where the model has instant access to the training data of all tasks, continual learning (CL) involves adapting to new sequentially arriving tasks over time without forgetting the previously acquired knowledge. Despite the wide practical adoption of CL and MTL and extensive literature on both areas, there remains a gap in the theoretical understanding of these methods when used with overparameterized models such as deep neural networks. This paper studies the overparameterized linear models as a proxy for more complex models. We develop theoretical results describing the effect of various system parameters on the model’s performance in an MTL setup. Specifically, we study the impact of model size, dataset size, and task similarity on the generalization error and knowledge transfer. Additionally, we present theoretical results to characterize the performance of replay-based CL models. Our results reveal the impact of buffer size and model capacity on the forgetting rate in a CL setup and help shed light on some of the state-of-the-art CL methods. Finally, through extensive empirical evaluations, we demonstrate that our theoretical findings are also applicable to deep neural networks, offering valuable guidance for designing MTL and CL models in practice  [^1] Code is available at .https://github.com/aminbana/MTL-Theory
@@ 1 Introduction
  • We provide explicit expressions describing the expected generalization error and knowledge transfer of multi-task learners and compare them to single-task learners in a linear regression setup with i.i.d. Gaussian features and additive noise. Our results help to understand the system’s behavior by highlighting the role and the interplay of various system parameters, including task similarity, model size, and sample size. We demonstrate that a peak exists in the generalization error curve of the multi-task learner when increasing model complexity. The error peak stems not only from label noise but also from the dissimilarity across the tasks. We also highlight the effect of each task’s sample size on the strength and location of the test error peak. Additionally, we measure knowledge transfer and indicate the conditions under which the tasks can be effectively learned together or interfere with each other.
  • We provide similar results for continual learners and use them to explain the characteristics of replay-based CL methods. These results demonstrate the impact of replay buffer size on forgetting and the generalization error. As a practical fruit, our results shed light on state-of-the-art replay-based CL methods by illustrating the effect of model size on the effectiveness of memory buffer, especially in the presence of practical limitations where the replay buffer and model size can not be arbitrarily large. Furthermore, we complete the theoretical results by analyzing the CL methods that use a combined strategy of regularization and rehearsal techniques.
  • By conducting experiments on practical datasets using DNNs, we empirically show that our findings for the linear models are generalizable to DNNs and can enlighten some of their characteristics. Specifically, we perform experiments using DNNs with varying model architectures on different practical datasets such as CIFAR-100, Imagenet-R, and CUB-200, to emphasize that the test error follows a similar trend observed in linear models, demonstrating that our results can help understand MTL in DNNs. We also perform experiments with replay-based deep continual learning methods and study the effect of memory buffer size in more depth.
@@ 2 Related Work
Additionally, some previous work studied multi-variate regression models, but their analysis or setup fundamentally differs from ours. For instance, they study sparse systems of equations (Lounici et al., 2009) with the goal of union support recovery (Kolar et al., 2011; Obozinski et al., 2011). As another example, Wu et al. (2020) establishes an upper bound on the error of regularized underparameterized models. As opposed to our work, none of these works are able to explicitly characterize the exact generalization performance of the system based on its fundamental components in the overparameterized regime.
@@
Most CL methods primarily use either model regularization (Schaul et al., 2015) or experience replay techniques (Kirkpatrick et al., 2017) to tackle catastrophic forgetting. Model regularization is based on consolidating the model parameters by penalizing drifts in the parameter space during model updates with the goal of preserving the older knowledge (Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018). The core idea for experience replay is to keep a portion of the training set in a separate memory buffer (Schaul et al., 2015). These samples are representative of previously encountered distributions and are replayed back during learning subsequent tasks along with the current task data. The stored samples enable the model to revisit and learn from the distributions it has encountered before, facilitating the retention of acquired knowledge from past tasks (Robins, 1995; Shin et al., 2017; Goodfellow et al., 2013). In practice, a combination of both techniques often leads to an optimal performance.
Several theoretical papers exist on understanding continual learning models (Doan et al., 2021; Yin et al., 2021; Chen et al., 2022). The most relevant study to our work is the investigation of regularization-based linear CL models. Goldfarb & Hand (2023) derive an upper bound for the generalization error of such models, and Lin et al. (2023) enhances this result by proposing a closed-form characterization of the error. However, these studies focus only on regularization-based techniques and are not helpful when studying replay-based CL methods. Besides, these results can be seen as a specific case of our studies in Section 4 by setting the memory size to zero.
@@
Double descent is not specific to DNNs (Belkin et al., 2019). For instance, linear regression exhibits similar behaviors under certain assumptions (Belkin et al., 2020). On the other hand, recent literature has pointed out a direct connection between linear models and more complex models such as neural networks optimized with SGD (Jacot et al., 2018; Gunasekar et al., 2018; Oymak & Soltanolkotabi, 2019). Consequently, several studies investigated linear models as a proxy for more complex models such as DNNs (Hastie et al., 2022). However, these works are primarily focused on single-task setups and are not capable of capturing task notions in MTL setups. In contrast, our work analyzes the MTL case where multiple tasks are learned together. Considering that even a simple multi-class classification problem is a special case of MTL, our work is essential in understanding commonly used overparameterized DNN models.
@@ 3.1 Problem Formulation
$y=\bar{x}_{t}^{T}\bar{w}_{t}^{*}+z,$ (1)
where $\bar{w}_{t}^{*}\in\mathbb{R}^{s_{t}}$ denotes the optimal parameters, and $z\in\mathbb{R}$ is the random noise (Lin et al., 2023; Belkin et al., 2018; Evron et al., 2022). The specific true feature space for each task is unknown, and the features for a specific task may not be useful to other tasks. Since we seek to use a single MTL model to learn all tasks, we consider a larger parameter space with size $p$ that encompasses all true features.
More formally, consider a global set of features indexed by $1,2,...$. We assume that the true set of features for task $t$ is denoted by $S_{t}$ such that $|S_{t}|=s_{t}$. Among all features, we chose a set of $p$ features denoted by $\mathcal{W}$ to train our model, assuming that $\cup_{t=1}^{T}S_{t}\subseteq\mathcal{W}$. With this in mind, we define the expanded ground truth for task $t$ by introducing $w_{t}^{*}\in\mathbb{R}^{p}$, where $w_{t}^{*}$ is the same as $\bar{w}_{t}^{*}$ in the $S_{t}$ indices and filled by zero in the remaining $p-s_{t}$ places.
@@
We consider a training dataset of size $n_{t}$ for task $t$ represented as $\mathcal{D}_{t}=\{(x_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, where $x_{t,i}\in\mathbb{R}^{p}$ and $y_{t,i}=x_{t,i}^{T}w_{t}^{*}+z_{t,i}$. We also assume the features and noise are i.i.d. according to the following distributions:
$x_{t,i}\sim\mathcal{N}(0,\mathbf{I}_{p})\quad\text{and}\quad z_{t,i}\sim\mathcal{N}(0,\sigma^{2}).$ (2)
@@
Consider a matrix representation by stacking the training data as $\mathbf{X}_{t}:=[x_{t,1},x_{t,2},...,x_{t,n_{t}}]\in\mathbb{R}^{p\times n_{t}}$, $\mathbf{y}_{t}:=[y_{t,1},y_{t,2},...,y_{t,n_{t}}]^{T}$, and the noise vector as $\mathbf{z}_{t}=[z_{t,1},z_{t,2},...,z_{t,n_{t}}]^{T}$, to summarize the data generation process as
$\mathbf{y}_{t}=\mathbf{X}_{t}^{T}w_{t}^{*}+\mathbf{z}_{t}.$ (3)
Similarly, we might stack the training data of all tasks to build $\mathbf{X}_{1T}\in\mathbb{R}^{p\times\bar{n}}$, $\mathbf{y}_{1T}\in\mathbb{R}^{\bar{n}}$ and $\mathbf{z}_{1T}\in\mathbb{R}^{\bar{n}}$ where $\bar{n}=\sum_{t=1}^{T}n_{t}$ is the total number of training samples.
@@
$w_{t}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\frac{1}{n_{t}}\|\mathbf{X}_{t}^{T}w-\mathbf{y}_{t}\|_{2}^{2}.$ (4)
In the underparameterized regime where $p<n_{t}$, there is a unique solution to minimizing the loss, given by $w_{t}=(\mathbf{X}_{t}\mathbf{X}_{t}^{T})^{-1}\mathbf{X}_{t}\mathbf{y}_{t}$. In contrast, in the overparameterized regime where $p>n_{t}$, there are infinite solutions with zero training error. Among all solutions, we are particularly interested in the solution with minimum $\ell_{2}$-norm, since it is the corresponding convergence point of applying stochastic gradient descent (SGD) on Equation 4 (Gunasekar et al., 2018). In fact, $w_{t}$ in this case is obtained by equivalently solving the following constrained optimization:
$w_{t}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w\|_{2}^{2}\quad\text{s.t.}\quad\mathbf{X}_{t}^{T}w=\mathbf{y}_{t}.$ (5)
@@
$\begin{split}\mathcal{L}_{t}(w)=\mathbb{E}_{x,z}[\|x^{T}w-y\|_{2}^{2}]=\sigma^{2}+\|w-w_{t}^{*}\|_{2}^{2}.\end{split}$ (6)
In what follows, we drop $\sigma^{2}$ and only use $\mathcal{L}_{t}(w)=\|w-w_{t}^{*}\|_{2}^{2}$ as a comparison criterion. Prior studies have characterized the behavior of the generalization error w.r.t. different model sizes in STL (Hastie et al., 2022). In fact, the error increases by increasing $p$ in the underparameterized regime until the training interpolation threshold $p\approx n_{t}$ is reached. After this point, the training error becomes zero and the model perfectly overfits the training noise. However, the model is benignly overfitted and increasing $p$ further will reduce the noise effect. This surprising behavior is associated with the implicit regularization of the SGD, as discussed in Section 2. A more detailed discussion of prior results is presented in Appendix A.
@@ 3.2 Main Results on Multi-Task Learning (MTL)
$w_{1T}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w\|_{2}^{2}\quad\text{s.t.}\quad\mathbf{X}_{1T}^{T}w=\mathbf{y}_{1T}.$ (7)
@@
    $K(w_{1T}):=\frac{1}{T}\sum_{t=1}^{T}\left[\mathcal{L}_{t}(w_{t})-\mathcal{L}_{t}(w_{1T})\right].$ (9)
@@
$\mathbb{E}[G(w_{1T})]=\underbrace{\sum_{t=1}^{T}\sum_{t^{\prime}=1}^{T}\frac{n_{t^{\prime}}}{Tp}(1+\frac{Tn_{t}}{2(p-\bar{n}-1)})\|w_{t}^{*}-w_{t^{\prime}}^{*}\|_{2}^{2}}_{\text{term $G_{1}$}}+\underbrace{\frac{1}{T}(1-\frac{\bar{n}}{p})\sum_{t=1}^{T}\|w_{t}^{*}\|_{2}^{2}}_{\text{term $G_{2}$}}+\underbrace{\frac{\bar{n}\sigma^{2}}{p-\bar{n}-1}}_{\text{term $G_{3}$}},$ (10)
$\begin{split}\mathbb{E}[K(w_{1T})]&=\underbrace{2\sum_{t=1}^{T}\sum_{t^{\prime}=1}^{T}\frac{n_{t^{\prime}}}{Tp}(1+\frac{Tn_{t}}{2(p-\bar{n}-1)})\langle w_{t}^{*},w_{t^{\prime}}^{*}\rangle}_{\text{term $K_{1}$}}\\ &-\underbrace{(1+\frac{1}{T}+\frac{\bar{n}}{p-\bar{n}-1})\sum_{t=1}^{T}\frac{n_{t}}{p}\|w_{t}^{*}\|_{2}^{2}}_{\text{term $K_{2}$}}-\underbrace{\frac{\bar{n}\sigma^{2}}{p-\bar{n}-1}+\frac{1}{T}\sum_{t=1}^{T}\frac{n_{t}\sigma^{2}}{p-n_{t}-1}}_{\text{term $K_{3}$}},\end{split}$ (11)
@@
where $\langle w_{t}^{*},w_{t^{\prime}}^{*}\rangle$ denotes the inner product of the two vectors and the expectation is due to the randomness of $X_{1T}$ and $z_{1T}$.
Due to the space limit, proofs for all theorems are provided in Appendix C. To the best of our knowledge, this is the first theoretical result on multi-task learning in overparameterized linear models that establishes a closed-form exact expression under non-asymptotic conditions, capturing the dependency of various system parameters. Setting $T=1$ recovers the prior STL results (Hastie et al., 2022), which means that our theorem is a more generalized version.
Theorem 3.1 precisely characterizes various phenomena that occur in the multi-task setting. Terms $G_{2}$ and $G_{3}$, which also appear in the STL setup, correspond to the task norm and noise strength. However, term $G_{1}$, which is specific to the MTL configuration, appears due to the distance between the optimal task vectors and is directly affected by task similarities. For a better demonstration, Figures 1(a) and 1(b) provide a visualization of the empirical average generalization error for different values of $p$, $\sigma$, and task similarities, as well as those predicted by our theory. Concretely, we have the the following observations:
@@
    Interpolation threshold shifts. Similar to the single-task learner, an interpolation threshold exists for the multi-task learner. For both models, the average error increases at the threshold and decreases afterward. However, the interpolation threshold for the single-task learner occurs at $p=n_{t}$, while for the multi-task learner, it happens at $p=\bar{n}$.
@@
    Descent floor exists when tasks are collaborative. When tasks are collaborative and the noise is not too strong, increasing the $p$ in the overparameterized regime can be a double-edged sword. On the one hand, increasing the number of parameters can help by reducing the effect of noise through the term $G_{3}$. On the other hand, increasing $p$ to $\infty$ can kill the positive knowledge transfer across the tasks. There exists a point in the middle where the effect of both mechanisms matches and a descent floor appears in the generalization error (see Figure 1(a) as an example).
@@
    Task dissimilarity and noise intensify the error peak. As previously mentioned, an error peak exists at the interpolation threshold. This behavior corresponds to the existence of the label noise and intensifies as $\sigma$ grows larger. However, in the single-task model, when $\sigma=0$, meaning that there exists no label noise in the training set, no error peak happens at the interpolation threshold (see Figures 1(a) and 1(b)). Meanwhile, there are two sources for the error peak in MTL. Not only can the noise cause the peak through the term $G_{3}$, but task dissimilarity can also contribute through the term $G_{1}$, meaning that highly dissimilar tasks can produce larger error peaks. However, the benign overfitting is still observable since both terms converge to zero as $p\to\infty$. This observation is consistent with earlier empirical findings where the double-descent phenomenon was observed even in the absence of label noise (Nakkiran et al., 2021).
Although these observations are derived from our theoretical analysis of linear models, in Section 5, we will use an empirical approach to demonstrate that DNNs trained with SGD also exhibit similar behaviors. Consequently, our theoretical insights can extend to and inform the understanding of more complex models.

### v2

@@ Abstract
Multi-task learning (MTL) is a machine learning paradigm that aims to improve the generalization performance of a model on multiple related tasks by training it simultaneously on those tasks. Unlike MTL, where the model has instant access to the training data of all tasks, continual learning (CL) involves adapting to new sequentially arriving tasks over time without forgetting the previously acquired knowledge. Despite the wide practical adoption of CL and MTL and extensive literature on both areas, there remains a gap in the theoretical understanding of these methods when used with overparameterized models such as deep neural networks. This paper studies the overparameterized linear models as a proxy for more complex models. We develop theoretical results describing the effect of various system parameters on the model’s performance in an MTL setup. Specifically, we study the impact of model size, dataset size, and task similarity on the generalization error and knowledge transfer. Additionally, we present theoretical results to characterize the performance of replay-based CL models. Our results reveal the impact of buffer size and model capacity on the forgetting rate in a CL setup and help shed light on some of the state-of-the-art CL methods. Finally, through extensive empirical evaluations, we demonstrate that our theoretical findings are also applicable to deep neural networks, offering valuable guidance for designing MTL and CL models in practice 111 Code is available at https://github.com/aminbana/MTL-Theory . .
@@ 1 Introduction
  • We provide explicit expressions describing the expected generalization error and knowledge transfer of multi-task learners and compare them to single-task learners in a linear regression setup with i.i.d. Gaussian features and additive noise. Our results help to understand the system’s behavior by highlighting the role and the interplay of various system parameters, including task similarity, model size, and sample size. For instance, we demonstrate that a peak exists in the generalization error curve of the multi-task learner when increasing model complexity and the error peak stems not only from label noise but also from the dissimilarity across the tasks. We also highlight the effect of the sample size on the strength and location of the test error peak. Additionally, we measure knowledge transfer and indicate the conditions under which the tasks can be effectively learned together or interfere with each other.
  • We provide similar results for continual learners and use them to explain the characteristics of replay-based CL methods. These results demonstrate the impact of replay buffer size on forgetting and the generalization error. Furthermore, we complete the theoretical results by analyzing the CL methods that use a combined strategy of relying on implicit regularization and rehearsal techniques. As a practical fruit, our results shed light on state-of-the-art replay-based CL methods by illustrating the effect of model size on reducing the effectiveness of memory buffer, especially in the presence of practical limitations where the replay buffer and model size can not be arbitrarily large.
  • We empirically show that our findings for the linear models are generalizable to DNNs by conducting various experiments using different architectures and practical datasets such as CIFAR-100, Imagenet-R, and CUB-200. Our experiments show that the test error follows a similar trend observed in linear models, indicating that our theoretical results can help understand MTL in DNNs. We also provide experiments with replay-based CL methods and study the effect of memory buffer size in practical setups with DNNs.
@@ 2 Related Work
Additionally, some previous work studied multi-variate regression models, but their analysis or setup fundamentally differs from ours. For instance, (Lounici et al., 2009) study sparse systems of equations with the goal of union support recovery (Kolar et al., 2011; Obozinski et al., 2011). More relevant to our work, Wu et al. (2020) analyzes a two-layer architecture with a shared module across tasks and separate output modules for each task. They derive an upper bound on the performance of an optimal multi-task learning (MTL) model in a two-task scenario, assuming data abundance and finite model width, and examine the impact of task alignment on MTL performance. In contrast to our work, these studies do not characterize the exact convergence point of SGD in the MTL setup or explicitly derive the system’s generalization performance in the overparameterized regime.
@@
CL methods predominantly use either model regularization (Kirkpatrick et al., 2017) or experience replay techniques (Schaul et al., 2015; Rolnick et al., 2019) to tackle catastrophic forgetting. Model regularization is based on consolidating the model parameters by penalizing drifts in the parameter space during model updates with the goal of preserving the older knowledge (Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018). The core idea for experience replay is to keep a portion of the training set in a separate memory buffer (Schaul et al., 2015). These samples are representative of previously encountered distributions and are replayed back during learning subsequent tasks along with the current task data. The stored samples enable the model to revisit and learn from the distributions it has encountered before, facilitating the retention of acquired knowledge from past tasks (Rolnick et al., 2019; Huang et al., 2024; Zhou et al., 2023). Memory-based methods have proven to be highly effective, and their variations often lead to the optimal approach in practice (Parisi et al., 2019).
Several theoretical papers exist on understanding continual learning models from different aspects (Doan et al., 2021; Yin et al., 2021; Chen et al., 2022). The most relevant to our work examines linear CL models under a Gaussian data model when trained sequentially using SGD as the data for new tasks arrive. For instance, Goldfarb & Hand (2023) derives an upper bound for the generalization error of such models, and Lin et al. (2023) refines this result by proposing a closed-form characterization of the error. Furthermore, Goldfarb et al. (2024) studies the interplay between forgetting and task similarity as a function of different levels of overparametrization under a more general data model. More recently, Guha & Lakshman (2024) extends the theoretical investigation to feed-forward neural networks to show that increasing network widths to reduce forgetting yields diminishing returns. Although these results are valuable and generally comply with our findings, they fall into the regularization CL-based methods. This is due to the fact that they merely rely on the implicit bias of SGD as a regularizer that tends to the solutions closer to the weights of the previous task, while in our study, we equip the continual learner with an external memory buffer and study the closed form characteristics of the error as a function of different system parameters, especially the memory size. Moreover, some of the mentioned results (e.g., Lin et al. (2023)) can be seen as a special case of our studies in Section 4 when memory size is zero.
@@
Double descent is not specific to DNNs (Belkin et al., 2019). For instance, linear regression exhibits similar behaviors under certain assumptions (Belkin et al., 2020). On the other hand, recent literature has pointed out a direct connection between linear models and more complex models such as neural networks optimized with SGD (Jacot et al., 2018; Gunasekar et al., 2018; Oymak & Soltanolkotabi, 2019). Consequently, several studies investigated linear models as a proxy for more complex models such as DNNs (Hastie et al., 2022). However, these works are primarily focused on single-task setups and are not capable of capturing task notions in MTL setups. In contrast, our work analyzes the MTL case where multiple tasks are learned together. Considering that even a simple multi-label classification problem is a special case of MTL, our work is essential in understanding commonly used overparameterized DNN models.
@@ 3.1 Problem Formulation
$y=\bar{x}_{t}^{\top}\bar{w}_{t}^{*}+z,$ (1)
where $\bar{w}_{t}^{*}\in\mathbb{R}^{s_{t}}$ denotes the optimal parameters, and $z\in\mathbb{R}$ is the random noise (Lin et al., 2023; Belkin et al., 2018). The specific true feature space for each task is unknown, and the features for a specific task may not be useful to other tasks. Since we seek to use a single MTL model to learn all tasks, we consider a larger parameter space with size $p$ that encompasses all true features.
More formally, consider a global set of features indexed by $1,2,\cdots$. We assume that the true set of features for task $t$ is denoted by $S_{t}$ such that $|S_{t}|=s_{t}$. Among all, we choose a set of $p$ features denoted by $\mathcal{W}$ to train our model, assuming that $\cup_{t=1}^{T}S_{t}\subseteq\mathcal{W}$. With this in mind, we define the expanded ground truth for task $t$ by introducing $w_{t}^{*}\in\mathbb{R}^{p}$, where $w_{t}^{*}$ is the same as $\bar{w}_{t}^{*}$ in the $S_{t}$ indices and filled by zero in the remaining $p-s_{t}$ places (Lin et al., 2023).
@@
We consider a training dataset of size $n_{t}$ for task $t$ represented as $\mathcal{D}_{t}=\{(x_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, where $x_{t,i}\in\mathbb{R}^{p}$ and $y_{t,i}=x_{t,i}^{\top}w_{t}^{*}+z_{t,i}$. We also assume the features and noise are i.i.d. according to the following distributions:
$x_{t,i}\sim\mathcal{N}(0,I_{p})\quad\text{and}\quad z_{t,i}\sim\mathcal{N}(0,\sigma^{2}).$ (2)
@@
Consider a matrix representation by stacking the training data as $X_{t}:=[x_{t,1},x_{t,2},...,x_{t,n_{t}}]\in\mathbb{R}^{p\times n_{t}}$, $y_{t}:=[y_{t,1},y_{t,2},...,y_{t,n_{t}}]^{\top}$, and the noise vector as $z_{t}=[z_{t,1},z_{t,2},...,z_{t,n_{t}}]^{\top}$, to summarize the data generation process as
$y_{t}=X_{t}^{\top}w_{t}^{*}+z_{t}.$ (3)
Similarly, we might stack the training data of all tasks to build $X_{1:T}\in\mathbb{R}^{p\times\bar{n}}$, $y_{1:T}\in\mathbb{R}^{\bar{n}}$ and $z_{1:T}\in\mathbb{R}^{\bar{n}}$ where $\bar{n}=\sum_{t=1}^{T}n_{t}$ is the total number of training samples.
@@
$w_{t}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\frac{1}{n_{t}}\|X_{t}^{\top}w-y_{t}\|_{2}^{2}.$ (4)
In the underparameterized regime where $p<n_{t}$, there is a unique solution to minimizing the loss, given by $w_{t}=(X_{t}X_{t}^{\top})^{-1}{X}_{t}y_{t}$. In contrast, in the overparameterized regime where $p>n_{t}$, there are infinite solutions with zero training error. Among all solutions, we are particularly interested in the solution with minimum $\ell_{2}$-norm, since it is the corresponding convergence point of applying stochastic gradient descent (SGD) on Equation 4 (Gunasekar et al., 2018) when starting from zero initialization. In fact, $w_{t}$ in this case is obtained by equivalently solving the following constrained optimization:
$w_{t}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w\|_{2}^{2}\quad\text{s.t.}\quad X_{t}^{\top}w=y_{t}.$ (5)
@@
$\begin{split}\mathcal{L}_{t}(w)=\mathbb{E}_{x,z}[(x^{\top}w-y)^{2}]=\|w-w_{t}^{*}\|_{2}^{2}+\sigma^{2}.\end{split}$ (6)
In what follows, we drop $\sigma^{2}$ and only use $\mathcal{L}_{t}(w)=\|w-w_{t}^{*}\|_{2}^{2}$ as a comparison criterion. As a prelude to our work, we review the following theoretical results on the generalization error of single-task learners:
Theorem 3.1.
$\mathbb{E}[\mathcal{L}_{t}(w_{t})]=\frac{p\sigma^{2}}{n_{t}-p-1},$ (7)
where the expectation is due to randomness of $X_{t}$ and $z_{t}$.
This result characterizes the generalization error of single-task learning (STL) across both the underparameterized and overparameterized regimes. To better understand this theorem, fix the number of training samples $n$ and analyze how the error evolves as a function of the number of parameters $p$. In the underparameterized regime ($p<n$) the learner doesn’t have enough capacity to fit all training samples. This serves as a regularization and leads to improved generalization, as the learned parameters approximate the optimal solution with low bias but some variance. As $p$ increases in this regime, the variance grows, leading to an overall increase in error until the interpolation threshold ($p\approx n$) is reached. At this critical point, there exists exactly one predictor that can perfectly fit the training data. However, this solution is highly sensitive to noise, as small fluctuations in the training data lead to large variations in the learned parameters, resulting in a sharp spike in test error.
Beyond the interpolation threshold, as $p$ continues to increase, the model enters the overparameterized regime, where it can completely memorize the training data, driving the training error to zero. Despite this apparent overfitting, the model exhibits benign overfitting, where increasing $p$ paradoxically reduces the influence of noise ($\frac{n_{t}\sigma^{2}}{p-n_{t}-1}\to 0$ as $p\to\infty$). As discussed in Section 2, this surprising behavior is associated with the implicit bias of SGD, which biases the optimization toward low-norm solutions. With this in mind, we are ready to present our results on multi-task learning.
@@ 3.2 Main Results on Multi-Task Learning (MTL)
$w_{1:T}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w\|_{2}^{2}\quad\text{s.t.}\quad X_{1:T}^{\top}w=y_{1:T}.$ (9)
@@
    $K(w_{1:T}):=\frac{1}{T}\sum_{t=1}^{T}\left[\mathcal{L}_{t}(w_{t})-\mathcal{L}_{t}(w_{1:T})\right].$ (11)
@@
$\mathbb{E}[G(w_{1:T})]=\underbrace{\sum_{t=1}^{T}\sum_{t^{\prime}=1}^{T}\frac{n_{t^{\prime}}}{Tp}(1+\frac{Tn_{t}}{2(p-\bar{n}-1)})\|w_{t}^{*}-w_{t^{\prime}}^{*}\|_{2}^{2}}_{\text{term $G_{1}$}}+\underbrace{\frac{1}{T}(1-\frac{\bar{n}}{p})\sum_{t=1}^{T}\|w_{t}^{*}\|_{2}^{2}}_{\text{term $G_{2}$}}+\underbrace{\frac{\bar{n}\sigma^{2}}{p-\bar{n}-1}}_{\text{term $G_{3}$}},$ (12)
$\begin{split}\mathbb{E}[K(w_{1:T})]&=\underbrace{2\sum_{t=1}^{T}\sum_{t^{\prime}=1}^{T}\frac{n_{t^{\prime}}}{Tp}(1+\frac{Tn_{t}}{2(p-\bar{n}-1)})\langle w_{t}^{*},w_{t^{\prime}}^{*}\rangle}_{\text{term $K_{1}$}}\\ &-\underbrace{(1+\frac{1}{T}+\frac{\bar{n}}{p-\bar{n}-1})\sum_{t=1}^{T}\frac{n_{t}}{p}\|w_{t}^{*}\|_{2}^{2}}_{\text{term $K_{2}$}}-\underbrace{\frac{\bar{n}\sigma^{2}}{p-\bar{n}-1}+\frac{1}{T}\sum_{t=1}^{T}\frac{n_{t}\sigma^{2}}{p-n_{t}-1}}_{\text{term $K_{3}$}},\end{split}$ (13)
where $\langle w_{t}^{*},w_{t^{\prime}}^{*}\rangle$ denotes the inner product of the two vectors and the expectation is due to the randomness of $X_{1:T}$ and $z_{1:T}$.
Proofs for all theorems are provided in Appendix A. To the best of our knowledge, this is the first theoretical analysis of multi-task learning in overparameterized linear models providing an exact closed-form expression under non-asymptotic conditions. Setting $T=1$ recovers the prior STL results (Hastie et al., 2022), which means that our theorem is a more generalized version of Theorem 3.1.
Theorem 3.2 precisely characterizes various phenomena that occur in the multi-task setting. Terms $G_{2}$ and $G_{3}$, which also appear in the STL setup, correspond to the task norm and noise strength. However, term $G_{1}$, which is specific to the MTL configuration, appears due to the distance between the optimal task vectors and is directly affected by task similarities. In fact, this observation motivates us to define a notion of similarity between any two tasks $t$ and $t^{\prime}$ as the Euclidean distance between the optimal task vectors, i.e. $\|w_{t}-w_{t^{\prime}}\|_{2}$.
@@
To achieve a better insight into the results of Theorem 3.2, Figures 1(a) and 1(b) provide a visualization of the empirical average generalization error for different values of $p$, $\sigma$, and task similarities, as well as those predicted by our theory. Concretely, we have the following observations:
@@
    Interpolation threshold shifts. Similar to the single-task learner, an interpolation threshold exists for the multi-task learner. For both models, the average error increases at the threshold and decreases afterward. However, the interpolation threshold for the single-task learner occurs at $p=n_{t}$, while for the multi-task learner, it happens at $p=\bar{n}$, which is the point where the number of MTL training samples matches the number of parameters.
@@
    Task dissimilarity and noise intensify the error peak. Recall that the STL test error peak that exists at the interpolation threshold appears because of the high variance of the learned predictor. This behavior stems only from the label noise and intensifies as $\sigma$ grows larger. Nevertheless, when $\sigma=0$, meaning that there exists no label noise in the training set, no error peak happens at the interpolation threshold of the single-task learner. Meanwhile, in an MTL setup, there are two sources for the error peak. Not only can the label noise cause the peak through the term $G_{3}$, but task dissimilarity can also contribute through the term $G_{1}$, meaning that highly dissimilar tasks can produce larger error peaks. In fact, the stochasticity of data points (i.e., $X$), together with high task dissimilarity, can lead to highly variant predictors with poor generalization behavior. This observation explains earlier empirical findings where the double-descent phenomenon was observed even in the absence of label noise (Nakkiran et al., 2021) and also distinguishes the study of MTL models from STL learners.
@@
    Descent floor exists when tasks are collaborative. When tasks are collaborative and the noise is not too strong, increasing $p$ in the overparameterized regime can be a double-edged sword. On one hand, increasing the number of parameters can help by reducing the effect of noise through the term $G_{3}$. On the other hand, increasing $p$ to $\infty$ can kill the positive knowledge transfer across the tasks. There exists a point in the middle where the effect of both mechanisms matches and a descent floor appears in the generalization error (see Figure 1(a) as an example).
Although these observations are derived from our theoretical analysis of linear models, in Section 5, we will use an empirical approach to demonstrate that DNNs trained with SGD also exhibit similar behaviors. Consequently, our theoretical insights can extend to and enhance the understanding of more complex models.